# OmniDraft: A Cross-vocabulary, Online Adaptive Drafter for On-device Speculative Decoding

### v2

@@ 4 Results
We refer to $\textnormal{SpD}_{DM}$ which is our implementation of SpD with direct mapping tokens only and is equivalent to UAG timor2025acceleratingllminferencelossless . N-gram postprocessing (pp) refers to using the N-gram cache as a postprocessing technique without directly training on it as we proposed in Algorithm 1. N-gram hit refers to the average number of successful N-gram cache lookups that were accepted by the target per speculative decoding step.
@@ 4.1 Cross-vocabulary Online Distillation
Target Method GSM8K MBPP+HumanEval Alpaca XSum Acc Rate Speedup Acc Rate Speedup Acc Rate Speedup Acc Rate Speedup Llama3-8B $\textnormal{SpD}_{DM}$ 0.10 0.94x 0.09 1.03x 0.09 0.96x 0.11 0.91x $\mathcal{L}_{\text{DM}}$ 0.32 1.58x 0.22 1.26x 0.16 1.25x 0.20 1.20x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ 0.42 1.70x 0.27 1.33x 0.20 1.30x 0.24 1.24x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ + LoRA 0.37 1.59x 0.19 1.28x 0.17 1.21x 0.23 1.21x Qwen2-7B $\textnormal{SpD}_{DM}$ 0.14 1.04x 0.09 0.91x 0.13 1.01x 0.12 0.96x $\mathcal{L}_{\text{DM}}$ 0.33 1.50x 0.22 1.29x 0.17 1.25x 0.19 1.16x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ 0.37 1.61x 0.26 1.36x 0.20 1.30x 0.22 1.22x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ + LoRA 0.31 1.41x 0.21 1.21x 0.18 1.25x 0.22 1.21x
@@ 4.3.1 Scalability to Larger Target and Draft Models
Target Method GSM8K MBPP+HumanEval Acc Rate Speedup Acc Rate Speedup Qwen2.5 7B $\textnormal{SpD}_{DM}$ (68M) 0.15 1.02x 0.10 0.94x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ (68M) 0.401 1.66x 0.27 1.33x $\textnormal{SpD}_{DM}$ (160M) 0.18 0.70x 0.13 0.60x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ (160M) 0.47 1.12x 0.32 0.90x Qwen2.5 14B $\textnormal{SpD}_{DM}$ (68M) 0.15 1.17x 0.10 1.14x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ (68M) 0.407 1.92x 0.272 1.57x $\textnormal{SpD}_{DM}$ (160M) 0.178 0.89x 0.13 0.84x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ (160M) 0.472 1.40x 0.33 1.19x Qwen2.5 32B $\textnormal{SpD}_{DM}$ (68M) 0.153 1.30x 0.10 1.23x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ (68M) 0.42 2.05x 0.274 1.71x $\textnormal{SpD}_{DM}$ (160M) 0.187 1.03x 0.133 0.97x $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ (160M) 0.49 1.62x 0.335 1.40x
@@ 4.3.3 Effectiveness of N-gram Cache
[…]
@@
[…]
@@ 4.3.4 Distillation Loss Comparisons
[…]
@@
[…]
@@ B.1 Speedup Metric
[…]
@@ B.5 Rank ablation for Cross-vocabulary online distillation of LoRA
We also ablate on the rank for the LoRA experiments, specifically focussed on the $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ + LoRA method. As shown in the Table 12, it is clear that as the rank increases, performance also improves, but beyond rank 32, there is a diminishing return in terms of overall improvement. As such we choose a conservative rank of 32 across all our experiments, tasks and methodologies. Moreover, for the “one drafter for all” setting, folding of the LoRA weights is infeasible and as such having a lower rank would benefit the overall memory and compute required for the drafter model on-device.
Table 12: Effect of rank for $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ + LoRA for Llama-68M and Llama3-8B on GSM8K
@@ Appendix C Additional Algorithm Details
1: Given target $p(\cdot)$, drafter $q_{\theta}(\cdot)$, online data stream $\mathcal{S}$, data buffer $\mathcal{Q}$, update interval $I$.
@@
1: Given target $p(\cdot)$, drafter $q_{\theta}(\cdot)$, acceptance prediction head $f_{\phi}(\cdot)$, online data stream $\mathcal{S}$, data buffer $\mathcal{Q}$, update interval $I$.
@@
1: Given target $p(\cdot)$, drafter $q_{\theta}(\cdot)$, acceptance prediction head $f_{\phi}(\cdot)$, online data stream $\mathcal{S}$, distillation data buffer $\mathcal{Q}$, adaptive drafting data buffer $\mathcal{R}$ with max size $N$, batch size $B$, update interval $I$.
@@ D.3 N-gram learning
From Figure 8, it can be seen that during the training of $\mathcal{L}_{\text{DM}}$ +$\lambda\mathcal{L}_{\text{N-gram}}$ on MBPP+HumanEval dataset, the average acceptance ratio for the n-grams is steadily improving, showing the impact of the loss function for better alignment of the n-grams. Furthermore, the impact is clearly visible even across different temperatures. Similarly as training progresses, the average number of successful cache hit for n-gram tokens per query is also improving.
@@ D.4 N-gram Cache Growth Rate
[…]

### v3

@@ 4 Results
We refer to $\textnormal{SpD}_{DM}$ as the baseline speculative decoding which uses direct mapping between vocabularies. N-gram postprocessing (pp) refers to using the N-gram cache as a postprocessing technique without directly training on it as we proposed in Algorithm 1. N-gram hit refers to the average number of successful N-gram cache lookups that were accepted by the target per speculative decoding step.
@@ 4.1 Cross-vocabulary Online Distillation
Target Method GSM8K MBPP+HumanEval Alpaca XSum Acc Rate Speedup Acc Rate Speedup Acc Rate Speedup Acc Rate Speedup Llama3-8B $\textnormal{SpD}_{DM}$ 0.10 0.94x 0.09 1.03x 0.09 0.96x 0.11 0.91x $\mathcal{L}_{\text{DM}}$ 0.32 1.58x 0.22 1.26x 0.16 1.25x 0.20 1.20x $\mathcal{L}_{\text{DM}}$ $\lambda\mathcal{L}_{\text{N-gram}}$ 0.42 1.70x 0.27 1.33x 0.20 1.30x 0.24 1.24x $\mathcal{L}_{\text{DM}}$ $\lambda\mathcal{L}_{\text{N-gram}}$ + LoRA 0.37 1.59x 0.19 1.28x 0.17 1.21x 0.23 1.21x Qwen2-7B $\textnormal{SpD}_{DM}$ 0.14 1.04x 0.09 0.91x 0.13 1.01x 0.12 0.96x $\mathcal{L}_{\text{DM}}$ 0.33 1.50x 0.22 1.29x 0.17 1.25x 0.19 1.16x $\mathcal{L}_{\text{DM}}$ $\lambda\mathcal{L}_{\text{N-gram}}$ 0.37 1.61x 0.26 1.36x 0.20 1.30x 0.22 1.22x $\mathcal{L}_{\text{DM}}$ $\lambda\mathcal{L}_{\text{N-gram}}$ + LoRA 0.31 1.41x 0.21 1.21x 0.18 1.25x 0.22 1.21x
@@ 4.3.1 Scalability to Larger Target and Draft Models
Target Method GSM8K MBPP+HumanEval Acc Rate Speedup Acc Rate Speedup Qwen2.5 7B $\textnormal{SpD}_{DM}$ (68M) 0.15 1.02x 0.10 0.94x $\mathcal{L}_{\text{DM}}$ $\lambda\mathcal{L}_{\text{N-gram}}$ (68M) 0.401 1.66x 0.27 1.33x $\textnormal{SpD}_{DM}$ (160M) 0.18 0.70x 0.13 0.60x $\mathcal{L}_{\text{DM}}$ $\lambda\mathcal{L}_{\text{N-gram}}$ (160M) 0.47 1.12x 0.32 0.90x Qwen2.5 14B $\textnormal{SpD}_{DM}$ (68M) 0.15 1.17x 0.10 1.14x $\mathcal{L}_{\text{DM}}$ $\lambda\mathcal{L}_{\text{N-gram}}$ (68M) 0.407 1.92x 0.272 1.57x $\textnormal{SpD}_{DM}$ (160M) 0.178 0.89x 0.13 0.84x $\mathcal{L}_{\text{DM}}$ $\lambda\mathcal{L}_{\text{N-gram}}$ (160M) 0.472 1.40x 0.33 1.19x Qwen2.5 32B $\textnormal{SpD}_{DM}$ (68M) 0.153 1.30x 0.10 1.23x $\mathcal{L}_{\text{DM}}$ $\lambda\mathcal{L}_{\text{N-gram}}$ (68M) 0.42 2.05x 0.274 1.71x $\textnormal{SpD}_{DM}$ (160M) 0.187 1.03x 0.133 0.97x $\mathcal{L}_{\text{DM}}$ $\lambda\mathcal{L}_{\text{N-gram}}$ (160M) 0.49 1.62x 0.335 1.40x
@@ 4.3.3 Effectiveness of N-gram Cache
In this section, we focus on the impact of the N-gram cache to different training variants as per Table 5. We perform ablation on a subset of GSM8K dataset (4k samples) across multiple techniques. It can be seen that the baseline $\textnormal{SpD}_{DM}$ performs poorly which indicates the mismatch between the drafter and the target model alignment for the pretrained models. We also capture all the possible N-gram matches during the $\textnormal{SpD}_{DM}$ baseline for the train set, and then use the cache as a lookup for an improved baseline in the $\textnormal{SpD}_{DM}$ + $\text{N-gram}_{pp}$. As such there is a cache hit of 0.87, which indicates that the baseline model without any pretraining can still benefit with the N-gram cache. We also train the model to improve the alignment using $\mathcal{L}_{\text{DM}}$ and although without the N-gram, we can still see a large improvement over the baseline. Moreover. when we train with $\textnormal{N-gram}_{pp}$, there is a further improvement on the overall speedup. Finally, we train using both $\mathcal{L}_{\text{DM}}$ + $\lambda\mathcal{L}_{\text{N-gram}}$, which provides the best results across all techniques on different draft length $k$.
@@
Metrics $\textnormal{SpD}_{DM}$ $\textnormal{SpD}_{DM}$ + $\text{N-gram}_{pp}$ $\mathcal{L}_{\text{DM}}$ $\mathcal{L}_{\text{DM}}$ + $\text{N-gram}_{pp}$ $\mathcal{L}_{\text{DM}}$ + $\lambda\mathcal{L}_{\text{N-gram}}$ $k=3$ Acc Rate 0.16 0.20 0.40 0.42 0.46 Speedup 1.04x 1.16x 1.59x 1.61x 1.66x Avg n-gram hit 0 0.87 0 0.48 2.40 $k=4$ Acc Rate 0.12 0.16 0.32 0.35 0.41 Speedup 1.01x 1.11x 1.51x 1.54x 1.61x Avg n-gram hit 0 1.49 0 0.75 3.66
@@ 4.3.4 Distillation Loss Comparisons
The different loss variants also provide different levels of performance on the test set as shown in Table 6. When trained only on the $\mathcal{L}_{\text{N-gram}}$, the training is very unstable. This could either be due to the number of n-grams being substantially smaller than the direct mapping tokens within most datasets, or the training requires additional constraints to direct towards a minima. Consequently, training on the combined loss provides better performance metrics across temperatures as well. The final $\mathcal{L}_{\text{DM}}$ KL + $\lambda\mathcal{L}_{\text{N-gram}}$ KL, 2, provides slightly better results on $\text{temperature}=1$ indicating the impact of the additional KL over the intersection vocabulary which is beneficial for the sampling process. However, we noticed tuning the scaling factor $\lambda$ becomes critical and hence we use Equation 1 for all of the experiments since it was much more stable across all tasks. We fix a $\lambda=0.2$ for all the tasks, across all experiments.
@@
Metrics $\mathcal{L}_{\text{N-gram}}$ NLL $\mathcal{L}_{\text{DM}}$ NLL + $\mathcal{L}_{\text{N-gram}}$ NLL $\mathcal{L}_{\text{DM}}$ KL $\mathcal{L}_{\text{N-gram}}$ NLL $\mathcal{L}_{\text{DM}}$ KL + $\mathcal{L}_{\text{N-gram}}$ KL $\text{Temperature}=0.01$ Acc Rate 0.090 0.368 0.375 0.376 Speedup 0.86x 1.51x 1.57x 1.57x $\text{Temperature}=1$ Acc Rate 0.070 0.271 0.273 0.275 Speedup 0.79x 1.23x 1.27x 1.31x
@@ B.1 Speedup Metric
We follow the convention in Medusa cai2024medusa to define the speedup metric. Given acceleration rate as the average number of tokens decoded per decoding step and overhead as the average per step latency of the proposed model divided by that of the vanilla model, speedup refers to the wall-time acceleration rate and can be computed with $\textbf{speedup}=\textbf{acceleration rate}/\textbf{overhead}$.
@@ B.5 Rank ablation for Cross-vocabulary online distillation of LoRA
We also ablate on the rank for the LoRA experiments, specifically focussed on the $\mathcal{L}_{\text{DM}}$ $\lambda\mathcal{L}_{\text{N-gram}}$ + LoRA method. As shown in the Table 12, it is clear that as the rank increases, performance also improves, but beyond rank 32, there is a diminishing return in terms of overall improvement. As such we choose a conservative rank of 32 across all our experiments, tasks and methodologies. Moreover, for the “one drafter for all” setting, folding of the LoRA weights is infeasible and as such having a lower rank would benefit the overall memory and compute required for the drafter model on-device.
Table 12: Effect of rank for $\mathcal{L}_{\text{DM}}$ $\lambda\mathcal{L}_{\text{N-gram}}$ + LoRA for Llama-68M and Llama3-8B on GSM8K
@@ Appendix C Additional Algorithm Details
1: Given target $p(\cdot)$, drafter $q_{\theta}(\cdot)$, online data stream $\mathcal{S}$, data buffer $\mathcal{Q}$, update interval II.
@@
1: Given target $p(\cdot)$, drafter $q_{\theta}(\cdot)$, acceptance prediction head $f_{\phi}(\cdot)$, online data stream $\mathcal{S}$, data buffer $\mathcal{Q}$, update interval II.
@@
1: Given target $p(\cdot)$, drafter $q_{\theta}(\cdot)$, acceptance prediction head $f_{\phi}(\cdot)$, online data stream $\mathcal{S}$, distillation data buffer $\mathcal{Q}$, adaptive drafting data buffer $\mathcal{R}$ with max size NN, batch size $B$, update interval II.
@@ D.3 N-gram learning
From Figure 8, it can be seen that during the training of $\mathcal{L}_{\text{DM}}$ $\lambda\mathcal{L}_{\text{N-gram}}$ on MBPP+HumanEval dataset, the average acceptance ratio for the n-grams is steadily improving, showing the impact of the loss function for better alignment of the n-grams. Furthermore, the impact is clearly visible even across different temperatures. Similarly as training progresses, the average number of successful cache hit for n-gram tokens per query is also improving.
@@ D.4 N-gram Cache Growth Rate
From Figure 9, We can define the growth rate of the N-gram cache to be $\text{growth\_rate}=\text{cache\_size}/\text{training\_tokens}$. From the tables, we can observe that the cache grows sub-linearly with number of tokens processed in online learning. The growth rate of the N-gram cache can also be seen to continually decrease.